OpenStudy (anonymous):

If the point (2, -2) is on the terminal side of the angle θ, find each of sin(θ),cos(θ),tan(θ) If the point (2, -2) is on the terminal side of the angle θ, find each of sin(θ),cos(θ),tan(θ) @Mathematics

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sin(\theta)=-\frac{\sqrt{2}}{2}\] etc

OpenStudy (anonymous):

do i set it like 2 on the x axis and -2 on the y-axis